# Improving Signed Propagation of Graph Neural Network Under Multiple Classes

### v6

@@ V-A2 Edge weight calibration for GAT
V-C Optimization
During the training phase, we apply the edge (Eq. 25) and confidence calibration (Eq. 36) to reduce the uncertainty as,
$\mathcal{L}_{total}=\mathcal{L}_{EGNN}+\lambda\mathcal{L}_{conf}$ (37)
Here, $\mathcal{L}_{EGNN}$ represents the node classification loss with the edge weight calibration, and $\lambda$ is a hyper-parameter that adjusts the intensity of the confidence calibration. As in Figure 3, the enhanced performance of signed GCN (with calibration) substantiates the significant performance improvement attributable to our methodology. Additionally, note that this strategy is agnostic to specific GNN methodologies, which enables broad applicability across other frameworks [10, 11, 9] since it leverages only node similarity and predictions.
(Reproducibility) Our code is in anonymous GitHub [^1] https://anonymous.4open.science/r/Signed-Calibrated-GNN-34B6/README.md. Also, we introduce the pseudo-code of applying our method to one of the signed GNNs, FAGCN [11], in Algorithm 1.
VI Experiments
We conduct extensive experiments to address the following research questions. We measure the node classification accuracy of all methods following the settings in [2], where the parameters are optimized solely with the training nodes.
Baselines. We employ several state-of-the-art methods for validation and comparison: (1) GNNs with positive edge weights: GCN [2], GAT [3], GIN [42], APPNP [43], GCNII [44], H2GCN [4], ACM-GCN [45], HOG-GCN [25], JacobiConv [46], and GloGNN [8]. (2) Edge pruning GNN: PTDNet [13]. (3) GNNs with signed propagation: GPRGNN [10], FAGCN [11], and GGCN [9].
VI-A Experimental Results (Q1)
In Table I, we describe the node classification accuracy of each method. We also show dissonance within parentheses. Dissonance is an uncertainty metric, a measure of effectiveness in distinguishing out-of-distribution data from conflict predictions [37, 47]:
$diss(\widehat{y}_{i})=\sum^{C}_{j=1}\left(\widehat{y}_{ij}\sum_{k\neq j}\widehat{y}_{ik}(1-{|\widehat{y}_{ik}-\widehat{y}_{ij}|\over\widehat{y}_{ij}+\widehat{y}_{ik}})\over\sum_{k\neq j}\widehat{y}_{ik}\right)$ (38)
The above equation can be computed for the non-zero values of $\widehat{y}_{i}$. The performance in Table I is inferior to the benchmarks since we chose the test scores with the best validation.
Homophily ratio plays an important role in GNNs. On the three homophilic citation networks, we observe that all methods perform well. As the homophily ratio decreases, methods that adjust weights based on associativity start to outperform plain GNNs. Similarly, methods of using signed messages (GPRGNN, FAGCN, and GGCN) or separating ego and neighbors (H2GCN) are shown to be effective in our experiments. However, for the two highly heterophilic graphs (Chameleon and Squirrel), where many nodes share the same neighborhoods [48], we notice that blocking information (PTDNet) outperforms signed GNNs.
Calibrations are effective in improving accuracy as well as in alleviating uncertainty. We apply calibrations ($\ddagger$) to three state-of-the-art methods that employ signed propagation (GPRGNN, FAGCN, and GGCN). The average improvements achieved by applying calibrations to these three methods are 4.37%, 3.1%, and 3.13%, respectively. Although calibrations help to improve the performances of plain GCN (2.65%), GAT (1.97%), and PTDNet (1.68%), it is shown to be more effective for signed messaging. In addition to this, we can also observe that the calibrated methods show lower dissonance (Eq. 38) compared to the corresponding vanilla model. To summarize, the results indicate that calibration not only reduces uncertainty but also improves accuracy significantly.
Figure 5: (Q2) Dissonance of vanilla GCN and its two variants; signed and zero-weight, which is the same as the one in Figure 3. The left one is Cora and the right one is the Chameleon dataset
VI-B Signed Propagation and Uncertainty (Q2)
To demonstrate the effect of signed messaging on uncertainty, we measure the dissonance using two variants of GCN; one variant assigns -1 and the other assigns zeros to heterophilic edges, respectively (please refer to §IV-C for details). The results are presented in Fig. 5 where the x-axis is the number of layers and the y-axis represents dissonance. Here, we conducted experiments with the Cora (Fig. 4a) and Chameleon (Fig. 4b) datasets. As stated in Thm. 4.6, the signed GCN variant in the green-colored line exhibits higher uncertainty compared to the zero-weight variant. In the Chameleon dataset, the entropy of vanilla GCN increases as the layer gets deeper, indicating that positive message-passing fails to deal adequately with heterophily. Although zero-weight GCN shows lower dissonance, signed GCN with calibrations consistently produces the smallest uncertainty, highlighting the efficacy of our method in mitigating such limitations.
Table III: (Q3) Ablation study. We describe the node classification accuracy with standard deviation by applying either the edge weight calibration or the confidence calibration to signed GNNs
  Datasets Cora Citeseer Actor Chameleon   GPRGNN 82.2 ± ${}_{\,\pm\,0.4}$ 70.4 ± ${}_{\,\pm\,0.8}$ 25.4 ± ${}_{\,\pm\,0.5}$ 49.1 ± ${}_{\,\pm\,0.9}$   + edge calib 83.3 ± ${}_{\,\pm\,0.6}$ 71.5 ± ${}_{\,\pm\,1.0}$ 26.3 ± ${}_{\,\pm\,0.6}$ 49.7 ± ${}_{\,\pm\,0.7}$   + conf calib 83.8 ± ${}_{\,\pm\,0.5}$ 72.6 ± ${}_{\,\pm\,0.5}$ 27.5 ± ${}_{\,\pm\,0.4}$ 50.2 ± ${}_{\,\pm\,0.3}$   FAGCN 81.9 ± ${}_{\,\pm\,0.5}$ 70.8 ± ${}_{\,\pm\,0.6}$ 25.2 ± ${}_{\,\pm\,0.8}$ 46.5 ± ${}_{\,\pm\,1.1}$   + edge calib 82.8 ± ${}_{\,\pm\,0.6}$ 72.3 ± ${}_{\,\pm\,0.7}$ 25.8 ± ${}_{\,\pm\,0.7}$ 46.9 ± ${}_{\,\pm\,1.0}$   + conf calib 83.5 ± ${}_{\,\pm\,0.3}$ 73.4 ± ${}_{\,\pm\,0.5}$ 26.3 ± ${}_{\,\pm\,0.4}$ 48.0 ± ${}_{\,\pm\,0.8}$   GGCN 81.0 ± ${}_{\,\pm\,1.2}$ 70.7 ± ${}_{\,\pm\,1.6}$ 22.5 ± ${}_{\,\pm\,0.5}$ 48.5 ± ${}_{\,\pm\,0.7}$   + edge calib 82.9 ± ${}_{\,\pm\,1.0}$ 71.0 ± ${}_{\,\pm\,1.1}$ 23.3 ± ${}_{\,\pm\,0.6}$ 48.9 ± ${}_{\,\pm\,0.7}$   + conf calib 83.7 ± ${}_{\,\pm\,0.9}$ 72.5 ± ${}_{\,\pm\,0.6}$ 23.6 ± ${}_{\,\pm\,0.4}$ 49.6 ± ${}_{\,\pm\,0.5}$
VI-C Ablation Study (Q3)
We conduct an ablation study to test the relative effectiveness of the two calibration methods. In Table III, we describe the node classification accuracy of signed GNNs on the homophilic (Cora, Citeseer) and heterophilic (Actor, Chameleon) graphs. Here, we apply either the edge weight calibration (+ edge calib) or the confidence calibration (+ conf calib). From the table, we can see that the methods equipped with the confidence calibration generally outperform the edge weight calibration and exhibit smaller variances. This is because the confidence calibration reduces the uncertainty of all nodes during training while the edge calibration is only applied to a (small) subset of edges during testing. Also, edge calibration is beneficial, preventing the propagation of signed information from similar nodes.
Figure 6: (Q4) Node classification accuracy w.r.t. the parameter $\lambda$ which controls the weight of the confidence calibration. Each figure represents a Chameleon (left) and Squirrel (right)
VI-D Parameter Sensitivity Analysis (Q4)
We conduct experiments to investigate the effect of the confidence calibration by adjusting $\lambda$ (Eq. 37). In Figure 6, we varied the values of $\lambda$ from 0 to 1. The edge weight calibration is combined with all methods along with the confidence calibration. The blue line represents PTDNet, while the others are signed GNNs. Notably, in both figures, calibrations improve the quality of all methods and are shown to be more effective in signed GNNs than in PTDNet, particularly in terms of relieving uncertainty. Also, finding an appropriate lambda is beneficial for overall improvement since assigning higher values to $\lambda$ generally degrades the overall performance, necessitating the use of early stopping during training. Additionally, please note that calibration can be limited by the inherent low capability of base models in heterophilic graphs, where it acts as a downstream task of GNNs.
VII Conclusion
We present new theoretical insights on the effect of using signed messages in multi-class graphs. In contrast to previous theorems that assume graphs with binary classes and solely focus on message-passing, we extend them to multi-class scenarios and introduce a new perspective on parameter updates. From this viewpoint, we highlight two critical limitations of using signed propagation: (1) It suffers from reduced separability in multi-class graphs with high possibility, and (2) It increases the probability of generating conflicting evidence compared to positive message-passing schemes. Based on these observations, we propose the edge weight and confidence calibrations as solutions to enhance performance and alleviate uncertainty. Our extensive performance experiments on six real-world datasets show that the calibration techniques are effective for both signed and plain GNNs. We believe that our theorems offer valuable insights for the development of improved aggregation schemes in future studies.

### v7

@@ F. Pseudo-code and Time Complexity of Our Algorithm
F-2. Time complexity of calibrated GCN
We analyze the computational complexity of calibrated GNN. For brevity, we take vanilla GCN (Kipf and Welling, 2016) as a base model. Generally, the cost of GCN is known to be proportional to the number of edges and trainable parameters $\mathcal{O}(|\mathcal{E}|\theta_{GCN})$. Here, $\theta_{GCN}$ is comprised of $\mathcal{O}(nz(X)F^{\prime}+F^{\prime}C)$ (Zhu et al., 2020), where $nz(\cdot)$ represents the non-zero elements of inputs and $F^{\prime}$ stand for the hidden dimension, and $C$ is the number of classes. Additionally, our method employs two types of calibration. The first one is edge weight calibration. For this, we need to retrieve the node features of each layer and calculate the cosine similarity for all connected nodes $|\mathcal{E}|^{2}$. Thus, the complexity becomes $\mathcal{O}(|\mathcal{E}|\theta_{GCN}+L|\mathcal{E}|^{2}$), where $L$ represents the number of entire layers. Additionally, confidence calibration takes $n=|\mathcal{V}_{valid}\cup\mathcal{V}_{test}|$ samples as inputs and finds top $k$ samples on each row of $\widehat{Y}$. Thus, their complexity can simply be defined as $\mathcal{O}(n+k)$. To summarize, the cost of calibrated GCN is $\mathcal{O}(2|\mathcal{E}|\theta_{GCN}+L|\mathcal{E}|^{2}+n+k)$, which is fairly efficient compared to plane algorithm.
G. Datasets and Implementations Details
G-1. Datasets
The statistical details of datasets are in Table 2. (1) Cora, Citeseer, Pubmed (Kipf and Welling, 2016) are citation graphs. where a node corresponds to a paper and edges are citations between them. The labels are the research topics of the papers. (2) Actor (Tang et al., 2009) is a co-occurrence graph of actors in the same movie. The labels are five types of actors. (3) Chameleon, Squirrel (Rozemberczki et al., 2019) are Wikipedia hyperlink networks. Each node is a web page and the edges are hyperlinks. Nodes are categorized into five classes based on monthly traffic.
G-2. General information
All methods including baselines and ours are implemented upon PyTorch Geometric [^2] https://pytorch-geometric.readthedocs.io/en/latest/modules/nn.html. For a fair comparison, we equalize the hidden dimension of the entire methodologies as 64. ReLU with dropout is used for non-linearity and to prevent over-fitting. We employ the log-Softmax as a cross-entropy function. The learning ratio is set to $1e^{-3}$ and the Adam optimizer is taken with weight decay $5e^{-4}$. For training, 20 nodes per class are randomly chosen and the remaining nodes are divided into two parts for validation and testing, where we followed the settings in (Kipf and Welling, 2016).
G-3. Implementation details about baseline methods
  • GCN (Kipf and Welling, 2016) is a first-order approximation of Chebyshev polynomials (Defferrard et al., 2016). For all datasets, we simply take 2 layers of GCN.
  • APPNP (Klicpera et al., 2018) combines personalized PageRank on GCN. We stack 10 layers and set the teleport probability ($\alpha$) as $\{0.1,0.1,0.1,0.5,0.2,0.3\}$ for Cora, Citeseer, Pubmed, Actor, Chameleon, and Squirrel.
  • GAT (Velickovic et al., 2017) calculates feature-based attention for edge coefficients. Similar to GCN, we construct 2 layers of GAT. The pair of (hidden dimension, head) is set as (8, 8) for the first layer, while the second layer is (1, # of classes).
  • GCNII (Chen et al., 2020) integrates an identity mapping function on APPNP. We set $\alpha=0.5$ and employ nine hidden layers. We increase the weight of identity mapping ($\beta$) that is inversely proportional to the heterophily of the dataset.
  • H2GCN (Zhu et al., 2020) suggests the separation of ego and neighbors during aggregation. We refer to the publicly available source code [^3] https://github.com/GemsLab/H2GCN for implementation.
  • PTDNet (Luo et al., 2021) removes disassortative edges before a message-passing. We also utilize the open source code [^4] https://github.com/flyingdoog/PTDNet and apply confidence calibration.
  • GPRGNN (Chien et al., 2020) generalized the personalized PageRank to deal with heterophily and over-smoothing. Referring to the open source code [^5] https://github.com/jianhao2016/GPRGNN, we tune the hyper-parameters based on the best validation score for each dataset.
  • FAGCN (Bo et al., 2021) determines the sign of edges using the node features. We implement the algorithm based on the sources [^6] https://github.com/bdy9527/FAGCN and tune the hyper-parameters concerning their accuracy.
  • GGCN (Yan et al., 2021) proposes the scaling of degrees and the separation of positive/negative adjacency matrices. We simply take the publicly available code [^7] https://github.com/Yujun-Yan/Heterophily_and_oversmoothing for evaluation.
Figure 9. By differentiating the number of classes, we compare the dissonance of signed GCN and zero-weight GCN using two graph variants. The leftmost value on the x-axis represents the original class, and the rightmost value represents the binary class
H. Further Evaluation
H-1. Case study
Through theoretical analyses, we confirm that signed messages increase the uncertainty in multi-class graphs. They have shown to be effective when $k^{\prime}$ gets closer to $-k$ (Eq. 18), but this probability is inversely proportional to the number of classes c. To further analyze this phenomenon, we compare the dissonance of two variants of GCN (signed GCN and zero-weight GCN) by decrement of the number of classes (c). Specifically, if the original data contains seven classes (e.g., Cora), we remove all the nodes that belong to the rightmost class to generate a graph with six labels. The results are illustrated in Figure 9. As can be seen, zero-weight GCN (red) tends to have lower dissonance under multiple classes. However, under binary classes (c=2), signed GCN (blue) shows lower uncertainty with the aid of ego-neighbor separation. In the binary case, zero-weight GCN only utilizes homophilous neighbors and fails to generalize under this condition.
Table 4. Node classification accuracy (%) on nine benchmark graphs. The gray-colored cell means the top 3 and bold with () indicates the best performances. Symbol $\ddagger$ means that calibration is applied to the baseline method
  Datasets Cora Citeseer Pubmed Actor Chameleon Squirrel Cornell Texas Wisconsin   $\mathcal{H}_{g}$ (Eq. 1) 0.81 0.74 0.8 0.22 0.23 0.22 0.11 0.06 0.16   MLP 53.2 ± ${}_{\,\pm\,0.5}$ 53.7 ± ${}_{\,\pm\,1.7}$ 69.7 ± ${}_{\,\pm\,0.4}$ 27.9 ± ${}_{\,\pm\,1.1}$ 41.2 ± ${}_{\,\pm\,1.8}$ 26.5 ± ${}_{\,\pm\,0.6}$ 60.1 ± ${}_{\,\pm\,1.2}$ 65.8 ± ${}_{\,\pm\,5.0}$ 73.5 ± ${}^{*}_{\,\pm\,5.4}$ GCN 79.0 ± ${}_{\,\pm\,0.6}$ 67.5 ± ${}_{\,\pm\,0.8}$ 77.6 ± ${}_{\,\pm\,0.2}$ 20.2 ± ${}_{\,\pm\,0.4}$ 49.3 ± ${}_{\,\pm\,0.5}$ 30.7 ± ${}_{\,\pm\,0.7}$ 39.4 ± ${}_{\,\pm\,4.3}$ 47.6 ± ${}_{\,\pm\,0.7}$ 40.5 ± ${}_{\,\pm\,1.9}$ GCN 81.0 ± ${}_{\,\pm\,0.9}$ 71.3 ± ${}_{\,\pm\,1.2}$ 77.8 ± ${}_{\,\pm\,0.4}$ 21.7 ± ${}_{\,\pm\,0.6}$ 49.4 ± ${}_{\,\pm\,0.6}$ 31.5 ± ${}_{\,\pm\,0.6}$ 39.6 ± ${}_{\,\pm\,3.1}$ 48.0 ± ${}_{\,\pm\,0.6}$ 40.5 ± ${}_{\,\pm\,1.4}$ GAT 80.1 ± ${}_{\,\pm\,0.6}$ 68.0 ± ${}_{\,\pm\,0.7}$ 78.0 ± ${}_{\,\pm\,0.4}$ 22.5 ± ${}_{\,\pm\,0.3}$ 47.9 ± ${}_{\,\pm\,0.8}$ 30.8 ± ${}_{\,\pm\,0.9}$ 42.1 ± ${}_{\,\pm\,3.1}$ 49.2 ± ${}_{\,\pm\,4.4}$ 45.8 ± ${}_{\,\pm\,5.3}$ GAT 81.4 ± ${}_{\,\pm\,0.4}$ 72.2 ± ${}_{\,\pm\,0.6}$ 78.3 ± ${}_{\,\pm\,0.3}$ 23.2 ± ${}_{\,\pm\,1.8}$ 49.2 ± ${}_{\,\pm\,0.4}$ 30.3 ± ${}_{\,\pm\,0.8}$ 42.5 ± ${}_{\,\pm\,2.5}$ 49.3 ± ${}_{\,\pm\,4.1}$ 47.1 ± ${}_{\,\pm\,5.0}$ GATv2 79.5 ± ${}_{\,\pm\,0.5}$ 67.4 ± ${}_{\,\pm\,0.6}$ 76.2 ± ${}_{\,\pm\,0.5}$ 22.1 ± ${}_{\,\pm\,2.0}$ 48.3 ± ${}_{\,\pm\,0.4}$ 28.9 ± ${}_{\,\pm\,1.2}$ 38.0 ± ${}_{\,\pm\,3.8}$ 52.5 ± ${}_{\,\pm\,1.7}$ 41.7 ± ${}_{\,\pm\,5.1}$ GIN 77.3 ± ${}_{\,\pm\,0.8}$ 66.1 ± ${}_{\,\pm\,0.6}$ 77.1 ± ${}_{\,\pm\,0.7}$ 24.6 ± ${}_{\,\pm\,0.8}$ 49.1 ± ${}_{\,\pm\,0.7}$ 28.4 ± ${}_{\,\pm\,2.2}$ 42.9 ± ${}_{\,\pm\,4.6}$ 53.5 ± ${}_{\,\pm\,3.0}$ 38.2 ± ${}_{\,\pm\,1.5}$ APPNP 81.3 ± ${}_{\,\pm\,0.5}$ 68.9 ± ${}_{\,\pm\,0.3}$ 79.0 ± ${}_{\,\pm\,0.3}$ 23.8 ± ${}_{\,\pm\,0.3}$ 48.0 ± ${}_{\,\pm\,0.7}$ 30.4 ± ${}_{\,\pm\,0.6}$ 49.8 ± ${}_{\,\pm\,3.6}$ 56.1 ± ${}_{\,\pm\,0.2}$ 45.7 ± ${}_{\,\pm\,1.7}$ GCNII 81.1 ± ${}_{\,\pm\,0.7}$ 68.5 ± ${}_{\,\pm\,1.4}$ 78.5 ± ${}_{\,\pm\,0.4}$ 25.9 ± ${}_{\,\pm\,1.2}$ 48.1 ± ${}_{\,\pm\,0.7}$ 29.1 ± ${}_{\,\pm\,0.9}$ 62.5 ± ${}_{\,\pm\,0.5}$ 69.3 ± ${}_{\,\pm\,2.1}$ 63.2 ± ${}_{\,\pm\,3.0}$ H2GCN 80.6 ± ${}_{\,\pm\,0.6}$ 68.2 ± ${}_{\,\pm\,0.7}$ 78.5 ± ${}_{\,\pm\,0.3}$ 25.6 ± ${}_{\,\pm\,1.0}$ 47.3 ± ${}_{\,\pm\,0.8}$ 31.3 ± ${}_{\,\pm\,0.7}$ 59.8 ± ${}_{\,\pm\,3.7}$ 66.3 ± ${}_{\,\pm\,4.6}$ 61.5 ± ${}_{\,\pm\,4.4}$ Ortho-GCN 80.6 ± ${}_{\,\pm\,0.4}$ 69.5 ± ${}_{\,\pm\,0.3}$ 76.9 ± ${}_{\,\pm\,0.3}$ 21.4 ± ${}_{\,\pm\,1.6}$ 46.7 ± ${}_{\,\pm\,0.5}$ 31.3 ± ${}_{\,\pm\,0.6}$ 45.4 ± ${}_{\,\pm\,4.7}$ 53.1 ± ${}_{\,\pm\,3.9}$ 46.6 ± ${}_{\,\pm\,5.8}$ P-reg 80.0 ± ${}_{\,\pm\,0.8}$ 69.2 ± ${}_{\,\pm\,0.7}$ 77.4 ± ${}_{\,\pm\,0.4}$ 20.9 ± ${}_{\,\pm\,0.5}$ 49.1 ± ${}_{\,\pm\,0.1}$ 33.6 ± ${}_{\,\pm\,0.4}$ 44.9 ± ${}_{\,\pm\,3.1}$ 58.5 ± ${}_{\,\pm\,4.2}$ 53.7 ± ${}_{\,\pm\,2.6}$ ACM-GCN 80.2 ± ${}_{\,\pm\,0.8}$ 68.3 ± ${}_{\,\pm\,1.1}$ 78.1 ± ${}_{\,\pm\,0.5}$ 24.9 ± ${}_{\,\pm\,2.0}$ 49.5 ± ${}_{\,\pm\,0.7}$ 31.6 ± ${}_{\,\pm\,0.4}$ 55.6 ± ${}_{\,\pm\,3.3}$ 58.9 ± ${}_{\,\pm\,2.6}$ 61.3 ± ${}_{\,\pm\,0.5}$ HOG-GCN 79.7 ± ${}_{\,\pm\,0.4}$ 68.2 ± ${}_{\,\pm\,0.6}$ 78.0 ± ${}_{\,\pm\,0.2}$ 21.5 ± ${}_{\,\pm\,0.5}$ 47.7 ± ${}_{\,\pm\,0.5}$ 30.1 ± ${}_{\,\pm\,0.4}$ 61.8 ± ${}_{\,\pm\,0.8}$ 60.4 ± ${}_{\,\pm\,0.6}$ 62.7 ± ${}_{\,\pm\,0.3}$ JacobiConv 81.9 ± ${}_{\,\pm\,0.6}$ 69.6 ± ${}_{\,\pm\,0.8}$ 78.5 ± ${}_{\,\pm\,0.4}$ 25.7 ± ${}_{\,\pm\,1.2}$ 52.8 ± ${}_{\,\pm\,0.9}$ 32.0 ± ${}_{\,\pm\,0.6}$ 55.3 ± ${}_{\,\pm\,3.4}$ 57.7 ± ${}_{\,\pm\,3.6}$ 53.4 ± ${}_{\,\pm\,1.6}$ GloGNN 82.4 ± ${}_{\,\pm\,0.3}$ 70.3 ± ${}_{\,\pm\,0.5}$ 79.3 ± ${}_{\,\pm\,0.2}$ 26.6 ± ${}_{\,\pm\,0.7}$ 48.2 ± ${}_{\,\pm\,0.3}$ 28.8 ± ${}_{\,\pm\,0.8}$ 56.8 ± ${}_{\,\pm\,1.1}$ 63.0 ± ${}_{\,\pm\,2.9}$ 59.5 ± ${}_{\,\pm\,1.3}$ TED-GCN 81.8 ± ${}_{\,\pm\,0.9}$ 71.4 ± ${}_{\,\pm\,0.6}$ 78.6 ± ${}_{\,\pm\,0.3}$ 26.0 ± ${}_{\,\pm\,0.9}$ 50.4 ± ${}_{\,\pm\,1.2}$ 33.0 ± ${}_{\,\pm\,0.9}$ 57.3 ± ${}_{\,\pm\,1.2}$ 60.5 ± ${}_{\,\pm\,1.6}$ 49.2 ± ${}_{\,\pm\,2.8}$ PCNet 81.5 ± ${}_{\,\pm\,0.8}$ 71.2 ± ${}_{\,\pm\,1.2}$ 78.8 ± ${}_{\,\pm\,0.3}$ 26.4 ± ${}_{\,\pm\,0.8}$ 48.1 ± ${}_{\,\pm\,1.7}$ 31.4 ± ${}_{\,\pm\,1.6}$ 55.0 ± ${}_{\,\pm\,1.7}$ 59.2 ± ${}_{\,\pm\,1.6}$ 54.4 ± ${}_{\,\pm\,1.1}$   PTDNet 81.2 ± ${}_{\,\pm\,0.9}$ 69.5 ± ${}_{\,\pm\,1.2}$ 78.8 ± ${}_{\,\pm\,0.5}$ 21.5 ± ${}_{\,\pm\,0.6}$ 50.6 ± ${}_{\,\pm\,0.9}$ 32.1 ± ${}_{\,\pm\,0.7}$ 63.3 ± ${}_{\,\pm\,1.1}$ 59.8 ± ${}_{\,\pm\,9.2}$ 53.1 ± ${}_{\,\pm\,5.6}$ PTDNet 81.9 ± ${}_{\,\pm\,0.6}$ 71.1 ± ${}_{\,\pm\,0.8}$ 79.0 ± ${}_{\,\pm\,0.2}$ 22.7 ± ${}_{\,\pm\,0.6}$ 50.9 ± ${}_{\,\pm\,0.3}$ 32.3 ± ${}_{\,\pm\,0.5}$ 64.4 ± ${}_{\,\pm\,3.3}$ 62.4 ± ${}_{\,\pm\,6.0}$ 55.7 ± ${}_{\,\pm\,3.7}$   GPR-GNN 82.2 ± ${}_{\,\pm\,0.4}$ 70.4 ± ${}_{\,\pm\,0.8}$ 79.1 ± ${}_{\,\pm\,0.1}$ 25.4 ± ${}_{\,\pm\,0.5}$ 49.1 ± ${}_{\,\pm\,0.7}$ 30.5 ± ${}_{\,\pm\,0.6}$ 57.1 ± ${}_{\,\pm\,1.6}$ 61.2 ± ${}_{\,\pm\,0.9}$ 62.4 ± ${}_{\,\pm\,1.2}$ GPRGNN 84.7 ± ${}^{*}_{\,\pm\,0.2}$ 73.3 ± ${}_{\,\pm\,0.5}$ 80.2 ± ${}^{*}_{\,\pm\,0.2}$ 28.1 ± ${}^{*}_{\,\pm\,1.3}$ 51.0 ± ${}_{\,\pm\,0.4}$ 31.8 ± ${}_{\,\pm\,0.4}$ 61.1 ± ${}_{\,\pm\,2.8}$ 63.0 ± ${}_{\,\pm\,1.4}$ 63.2 ± ${}_{\,\pm\,1.5}$ FAGCN 81.9 ± ${}_{\,\pm\,0.5}$ 70.8 ± ${}_{\,\pm\,0.6}$ 79.0 ± ${}_{\,\pm\,0.5}$ 25.2 ± ${}_{\,\pm\,0.8}$ 46.5 ± ${}_{\,\pm\,1.1}$ 30.4 ± ${}_{\,\pm\,0.4}$ 55.6 ± ${}_{\,\pm\,3.0}$ 50.1 ± ${}_{\,\pm\,3.9}$ 57.8 ± ${}_{\,\pm\,6.5}$ FAGCN 84.1 ± ${}_{\,\pm\,0.4}$ 73.8 ± ${}^{*}_{\,\pm\,0.5}$ 79.7 ± ${}_{\,\pm\,0.2}$ 27.6 ± ${}_{\,\pm\,0.5}$ 48.8 ± ${}_{\,\pm\,0.7}$ 31.3 ± ${}_{\,\pm\,0.5}$ 60.5 ± ${}_{\,\pm\,2.9}$ 55.4 ± ${}_{\,\pm\,2.6}$ 59.6 ± ${}_{\,\pm\,4.6}$ GGCN 81.0 ± ${}_{\,\pm\,1.2}$ 70.7 ± ${}_{\,\pm\,1.6}$ 78.2 ± ${}_{\,\pm\,0.4}$ 22.5 ± ${}_{\,\pm\,0.5}$ 48.5 ± ${}_{\,\pm\,0.7}$ 30.2 ± ${}_{\,\pm\,0.6}$ 62.8 ± ${}_{\,\pm\,4.0}$ 66.4 ± ${}_{\,\pm\,1.7}$ 61.7 ± ${}_{\,\pm\,3.2}$ GGCN 83.9 ± ${}_{\,\pm\,0.8}$ 73.0 ± ${}_{\,\pm\,0.4}$ 78.9 ± ${}_{\,\pm\,0.3}$ 24.6 ± ${}_{\,\pm\,0.4}$ 50.0 ± ${}_{\,\pm\,0.4}$ 31.1 ± ${}_{\,\pm\,0.6}$ 66.7 ± ${}^{*}_{\,\pm\,2.2}$ 69.9 ± ${}^{*}_{\,\pm\,0.6}$ 62.6 ± ${}_{\,\pm\,1.5}$
H-2. Node classification accuracy of other state-of-the-art methods
In Table 4, we measure the node classification accuracy by adding three datasets and various state-of-the-art baselines. The WebKB graphs Cornell, Texas, Wisconsin [^8] http://www.cs.cmu.edu/ webkb/ are comprised of the web pages of computer science departments in various universities. We briefly introduce the details of newly added methods and their implementation as follows:
  • P-reg (Yang et al., 2021) ensembles a regularization term to provide additional information that training nodes might not capture.(source code [^9] https://github.com/yang-han/P-reg)
  • ACM-GCN (Luan et al., 2022) suggests a local diversification operation through the adaptive channel mixing algorithm. (source code [^10] https://github.com/SitaoLuan/ACM-GNN)
  • HOG-GCN (Wang et al., 2022) adaptively controls the propagation mechanism by measuring the homophily degrees between two nodes. (source code [^11] https://github.com/hedongxiao-tju/HOG-GCN)
  • JacobiConv (Wang and Zhang, 2022) studies the expressive power of spectral GNN and establishes a connection with the graph isomorphism testing. (source code [^12] https://github.com/GraphPKU/JacobiConv)
  • GloGNN (Li et al., 2022) receives information from global nodes, which can accelerate neighborhood aggregation. (source code [^13] https://github.com/RecklessRonan/GloGNN)
  • TED-GCN (Yan et al., 2024) redefines GCN’s depth $L$ as a trainable parameter, which can control its signal processing capability to model both homophily/heterophily graphs.
  • PCNet (Li et al., 2024) proposes a two-fold filtering mechanism to extract homophily in heterophilic graphs (source code) [^14] https://github.com/uestclbh/PC-Conv.
In the above table, we can see that signed GNNs (GPRGNN, FAGCN, and GGCN) with calibration generally outperform the baseline methods. As mentioned before, signed methods achieve inferior performance on the Chameleon and Squirrel datasets, where many nodes share the same neighboring nodes. Under these conditions, PTDNet outperforms the signed methods, which justifies our analysis. However, in another graph, calibrated($\ddagger$) GNNs still achieve competitive results. Especially, for the WebKB datasets, the average improvement of signed GNNs through calibration is 3.7%, 7.5%, and 4.4%, respectively. This result demonstrates that calibration can significantly enhance signed propagation, leading to state-of-the-art performance on real-world benchmark graph datasets.
Remark. Note that the edge classification error ratio ($e$) in Equation 18 is proportional to the performance of the GNN used as the base model. Thus, if the value of $e$ for a model is high enough, blocking information might help improve accuracy rather than the signed propagation.